Number of Credits / Fees
3 / 2000
6 / 2600
9 / 3200
12 / 3800
15 / 4400
(a) Write an equation to calculate the fees F when the number of credits n is given?

(b) Use this equation from part (a) to calculate the fee for 17 credits.



Answer by stanbon(75887)

Number of Credits / Fees
3 / 2000
6 / 2600
9 / 3200
12 / 3800
15 / 4400
(a) Write an equation to calculate the fees F when the number of credits n is given?
You can see that for every increase of 3 credits the fees increases by 600.
So the function is linear.
----------------------------------------
Use any two points to find the linear equation:
slope = (2600-2000)/(6-3) = 600/3 = 200
Form y = mx + b
2000 = 200*3 + b
b = 1400
Equation: fee(x) = 200x + 1400 where x is the number of credits
---------------------------------------------------------------------
(b) Use this equation from part (a) to calculate the fee for 17 credits.
fee(17) = 200*17+1400 = $4800
===============================
Cheers,
Stan H.